OpenStudy (anonymous):

2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

What do you want solved with these two equations?

OpenStudy (anonymous):

@hottgurl ?

OpenStudy (anonymous):

Solve the system of linear equations @jomary

OpenStudy (anonymous):

immm bout to send u the answers they have...@jomary

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a.x = 0, y = 2 3 b.x = 2, y = - 2 3 c.x = 4 3 , y = 2 9 d.x = 4, y = -2

OpenStudy (anonymous):

Mkay Its A. x=0 and y = 2/3

OpenStudy (anonymous):

okay thank you so much @jomary can u help me so more please

OpenStudy (anonymous):

Your welcome! Sure thing, fire away! :D

OpenStudy (anonymous):

Solve the system by substitution. 6x + 3y = -6 2x + y = -2

OpenStudy (anonymous):

Are there answer choices?

OpenStudy (anonymous):

x = 0, y = -2 infinite solutions x = -1, y = 0 no solution

OpenStudy (anonymous):

@jomary

OpenStudy (anonymous):

I'm here...just working. one moment

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x = -1, y = 0. =3

OpenStudy (anonymous):

thank you very much.